# HyperMARL: Adaptive Hypernetworks for Multi-Agent RL

### v2

@@ A.1 Quantifying Team Diversity
In contrast to  (Bettini et al., 2023), we use Jensen-Shannon Distance (JSD) (Endres & Schindelin, 2003; Lin, 1991) as $D$, rather than the Wasserstein metric (Vaserstein, 1969). As shown in Appendix A.2, JSD is a robust metric for both continuous and discrete cases, and provides a more reliable measure of policy distance.
A.2 Finding a Suitable Distance Function for Policy Diversity
@@
(Bettini et al., 2024) use the Wasserstein metric for continuous policies (Vaserstein, 1969) as distance function $D$, while (McKee et al., 2022) use the total variation distance for discrete policies. For discrete policies, Wasserstein distance would require a cost function representing the cost of changing from one action to another, which might not be trivial to come up with. On the other hand, although well-suited for discrete policies, TVD might miss changes in action probabilities because it measures the largest difference assigned to an event (i.e. action) between two probability distributions.
We consider a simple example to illustrate this point. Suppose we have two policies $\pi^{1}$ and $\pi^{2}$ with action probabilities as shown in Figure 8. $\pi^{1}$ stays constant, while $\pi^{2}$ changes gradually over timesteps. We see that even as $\pi^{2}$ changes over time, the $TVD(\pi^{1},\pi^{2})$ between $\pi^{1}$ and $\pi^{2}$ remains constant. This is because TVD only measures the largest difference between the two distributions, and does not consider the overall difference between them. On the other hand, the Jensen-Shannon distance (JSD) (Endres & Schindelin, 2003), which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function. Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetry, and it satisfies the triangle inequality.
@@
Jensen-Shannon Divergence (Lin, 1991) Both Divergence Yes $JSD(p\parallel q)=\frac{1}{2}D_{KL}(p\parallel m)+\frac{1}{2}D_{KL}(q\parallel m),\,m=\frac{1}{2}(p+q)$
Jensen-Shannon Distance (Endres & Schindelin, 2003) Both Metric Yes $\sqrt{JSD(p\parallel q)}$
@@
Appendix B Specialisation and Synchronisation Game
Theorem B.1.
A shared policy cannot learn the optimal behaviour for the two-player Specialisation Game.
@@ Appendix B Specialisation and Synchronisation Game
In Table 6, we present the results for the Specialisation and Synchronisation Games when actions are sampled during evaluation rather than selecting the argmax, as was done in Table 2.
Table 6: Average Evaluation Reward at the end of training (mean $\pm$ std) for REINFORCE variants in the Specialisation and Synchronisation Games, across 2, 4, and 8 Agents. Bold indicates values within 1 std of the maximum. Actions are sampled during evaluation.
Specialisation Game
Method 2 4 8
NoPS .9996 $\pm$ .0004 .9988 $\pm$ .0007 .9989 $\pm$ .0008
FuPS .7493 $\pm$ .0062 .6826 $\pm$ .0039 .6545 $\pm$ .0025
FuPS+ID .$9$$\pm$ .0002 .9995 $\pm$ .0003 .7275 $\pm$ .0293
Synchronisation Game
Method 2 4 8
NoPS .9993 $\pm$ .0004 .9976 $\pm$ .0022 .1250 $\pm$ 0
FuPS .$9$$\pm$ .0002 1.0 $\pm$ 0 .3000 $\pm$ .3500
FuPS+ID .$9$$\pm$ 0 .$9$$\pm$ .0003 .1250 $\pm$ 0
Appendix C HyperMARL Details
C.1 HyperMARL Pseudocode
@@
4:   Agent embeddings $\{e^{i}\}_{i=1}^{n}$One-hot or orthogonally initialized learnable parameters}
@@
C.2 Scalability and Parameter Efficiency
@@
To reduce parameter count, techniques like shared hypernetworks, chunked hypernetworks (von Oswald et al., 2020; Chauhan et al., 2024), or producing low-rank weight approximations, can be used. While naive implementations are parameter-intensive, this might be less critical in RL and MARL which commonly have smaller actor-critic networks. Moreover, HyperMARL’s near-constant scaling with agents suggests strong potential for large-scale MARL applications.
@@
Appendix D Experiment Details
D.1 Environments
@@
Dispersion (VMAS) (Bettini et al., 2022): A 2D environment where four agents collect unique food particles. This task requires specialised heterogeneous behaviours and resembles the Specialisation Game from Section 3.1.
@@
Navigation (VMAS) (Bettini et al., 2022): Agents navigate in a 2D space to assigned goals, receiving dense rewards based on proximity. Configurations include shared goals (homogeneous), unique goals (heterogeneous), and mixed goals, where some agents share goals while others have unique ones.
@@
Multi-Agent MuJoCo (MAMuJoCo) (Peng et al., 2021): A multi-agent extension of MuJoCo, where robot body parts (e.g., a cheetah’s legs) are modelled as different agents. Agents coordinate to perform efficient motion, requiring heterogeneous policies (Zhong et al., 2024).
@@
SMAX (JaxMARL) (Rutherford et al., 2024): Discrete action tasks with 2 to 20 agents on SMACv1- and SMACv2-style maps. FuPS baselines have been shown optimal for these settings (Yu et al., 2022; Fu et al., 2022) indicating homogeneous behaviour is preferred here.
Table 7: Overview of Baseline Methods and Justifications. Selected baselines (✓) are chosen based on their relevance to our focus on parameter sharing and specialisation/generalisation in MARL. Non-selected methods (✗) are excluded due to implementation limitations or misalignment with our scope.
Baseline Method Category Chosen Justification
IPPO (NoPS, FuPS) (De Witt et al., 2020) NoPS/FuPS Serves as a strong MARL baseline for both independent (NoPS) and shared (FuPS) policies.
MAPPO (NoPS, FuPS) (Yu et al., 2022) NoPS/FuPS Provides a robust MARL baseline with centralized critics for both NoPS and FuPS configurations.
DiCo (Bettini et al., 2024) Architectural Enables comparison with a method that employs preset diversity levels and maintains both non-shared and shared parameters.
HAPPO (Zhong et al., 2024) Sequential Updates Facilitates comparison with a method designed for heterogeneous behaviours, utilizing sequential updates and maintaining non-shared parameters per agent.
@@
SEAC (Christianos et al., 2020) Shared Experience Focuses on experience sharing rather than parameter sharing, which is outside our scope.
@@
SePS (Christianos et al., 2021) SePS Requires pretraining, which is beyond the scope of our current study.
@@
CDAS (Li et al., 2021) Intrinsic Reward Only implemented for off-policy methods which are outside our scope, and has been shown to underperform FuPS/NoPS (Fu et al., 2022).
ROMA/RODE (Wang et al., 2020b, a) Role-based Demonstrates limited practical performance (Christianos et al., 2020).
SNP-PS (Kim & Sung, 2023) Architectural/Pruning No available code implementation.
Kaleidoscope (Li et al., 2024) Architectural/Pruning Only implemented for off-policy methods, which is outside our scope.
D.1.1 Training and Evaluation

### v3

@@ D.1 Quantifying Team Diversity
In contrast to  [7], we use Jensen-Shannon Distance (JSD) [13, 27] as $D$, rather than the Wasserstein metric [41]. As shown in Appendix D.2, JSD is a robust metric for both continuous and discrete cases, and provides a more reliable measure of policy distance.
D.2 Finding a Suitable Distance Function for Policy Diversity
@@
[6] use the Wasserstein metric for continuous policies [41] as distance function $D$, while [29] use the total variation distance for discrete policies. For discrete policies, Wasserstein distance would require a cost function representing the cost of changing from one action to another, which might not be trivial to come up with. On the other hand, although well-suited for discrete policies, TVD might miss changes in action probabilities because it measures the largest difference assigned to an event (i.e. action) between two probability distributions.
We consider a simple example to illustrate this point. Suppose we have two policies $\pi^{1}$ and $\pi^{2}$ with action probabilities as shown in Figure 10. $\pi^{1}$ stays constant, while $\pi^{2}$ changes gradually over timesteps. We see that even as $\pi^{2}$ changes over time, the $TVD(\pi^{1},\pi^{2})$ between $\pi^{1}$ and $\pi^{2}$ remains constant. This is because TVD only measures the largest difference between the two distributions, and does not consider the overall difference between them. On the other hand, the Jensen-Shannon distance (JSD) [13], which is the square root of the Jensen-Shannon divergence, does not have this problem as it is a smooth distance function. Furthermore, it satisfies the conditions for being a metric – it is non-negative, symmetry, and it satisfies the triangle inequality.
@@
Jensen-Shannon Divergence [27] Both Divergence Yes $JSD(p\parallel q)=\frac{1}{2}D_{KL}(p\parallel m)+\frac{1}{2}D_{KL}(q\parallel m),\,m=\frac{1}{2}(p+q)$
Jensen-Shannon Distance [13] Both Metric Yes $\sqrt{JSD(p\parallel q)}$
@@
Appendix E Specialisation and Synchronisation Games
To study the challenges of specialisation and coordination in an isolated setting, we introduce the Specialisation and Synchronisation Games, drawing inspiration from a version of the XOR game [16], VMAS’s Dispersion [5] and coordination and anti-coordination games in game theory [32]. These environments encourage agents to take distinct actions (Specialisation) or take identical actions (Synchronisation). Despite their deceptively simple payoff structure, these games present substantial learning challenges – non-stationary reward distributions driven by others’ adapting behaviours and in their temporal extension, the joint observation spaces grows exponentially with the number of agents.
E.1 Specialisation and Synchronisation Games Description
Specialisation Game. Agents are encouraged to choose distinct actions. In the simplest setting, it is a two-player matrix game where each agent selects between two actions ($A$ or $B$). As shown in Figure 2(a), agents receive a payoff of $1.0$ when their actions differ (creating two pure Nash equilibria on the anti-diagonal) and $0.5$ when they match. This structure satisfies Definition 1, since optimal joint policies require complementary, non-identical strategies. There is also a symmetric mixed-strategy equilibrium in which each agent plays $A$ and $B$ with probability $0.5$.
Synchronisation Game. Conversely, agents are encouraged to coordinate and choose identical actions. The payoff matrix inverts the Specialisation game’s structure, agents receive $1$ for matching actions and $0.5$ for differing ones. This creates two pure Nash equilibria along the diagonal of the payoff matrix (Figure 2(b)), and incentivises uniform behaviour across agents.
$N$-Agent Extension. Both games naturally scale to $n$ agents and $n$ possible actions. In Specialisation, unique actions receive a payoff of $1.0$, while selecting the same action receives payoffs of $\frac{1}{k}$, where $k$ is the number of agents choosing that action. In contrast, in Synchronisation, agents receive maximum payoffs ($1.0$) only when all actions match. For partial coordination, rewards follow a hyperbolic scale, $\frac{1}{n-k+1}$, encouraging agents to align their choices. Visualisations in Figure 2 and detailed reward profiles appear in Figure 12.
E.2 General-$n$ Payoff Definitions
Both games generalise naturally to $n$ agents and $n$ possible actions. We show the reward profiles for $n=5$ agents in Figure 12.
Let $\mathbf{a}=(a^{1},\dots,a^{n})\in\{1,\dots,n\}^{n}$ and $k_{a}=\bigl{|}\{\,j:a^{j}=a\}\bigr{|}$ be the joint action profile and the count of agents choosing action $a$, respectively.
Temporal Extension. To model sequential decision-making, we extend each normal-form game into a repeated Markov game, where the state at time $t$ is the joint action at time $t-1$. At each step $t$ all agents observe $a^{t-1}$, choose $a_{i}^{t}$, and receive the original Specialisation or Synchronisation payoff. This repeated setup isolates how agents adapt based on past joint actions, exposing temporal patterns of specialisation and coordination.
E.2.1 Specialisation Game
The reward is formulated as follows:
$r_{\mathrm{spec}}^{\,i}(\mathbf{a})=\begin{cases}1.0,&\text{if }k_{a^{\,i}}=1\quad\bigl{(}\text{unique action}\bigr{)};\\[6.0pt] \displaystyle\frac{1}{k_{a^{\,i}}},&\text{if }k_{a^{\,i}}>1\quad\bigl{(}\text{shared action}\bigr{)}.\end{cases}$
E.2.2 Synchronisation Game
The reward is formulated as follows:
$r_{\mathrm{sync}}^{\,i}(\mathbf{a})=\frac{1}{\,n-k_{a^{\,i}}+1\,},$
so that $r_{\mathrm{sync}}^{\,i}=1.0$ when $k_{a^{\,i}}=n$ (all agents select the same action), and otherwise follows a hyperbolic scale encouraging consensus.
Figure 12: Reward profiles for the Specialisation (blue) and Synchronisation (orange) games with $n=5$ agents. In the Specialisation game, an agent’s payoff peaks when it selects a unique action, and then decays as when actions are shared. In the Synchronisation game, payoffs follow a hyperbolic scale $1/(n-k+1)$, reaching maximum only under full consensus, thereby incentivising alignment.
E.3 Proof that FuPS cannot represent the optimal policy in the two-player Specialisation Game
Theorem 1.
A stochastic, shared policy without agent IDs cannot learn the optimal behaviour for the two-player Specialisation Game.
@@ E.3 Proof that FuPS cannot represent the optimal policy in the two-player Specialisation Game
E.4 Measuring Agent Gradient Conflict
We measure gradient conflicts, via the cosine similarity between agents’ gradients $\cos\bigl{(}g_{t}^{(i)},g_{t}^{(j)}\bigr{)}=\frac{\langle g_{t}^{(i)},g_{t}^{(j)}\rangle}{\|g_{t}^{(i)}\|\|g_{t}^{(j)}\|}$, where $g_{t}^{(i)}=\nabla_{\theta}\mathcal{L}^{(i)}(\theta_{t})$.
E.5 Empirical Results in N-player Specialisation and Synchronisation Normal-Form Game
To assess this limitations of FuPS and NoPS in practice, we compare three REINFORCE [46] variants in both games with $n=2,4,8,16,32$ agents: NoPS (No Parameter Sharing), FuPS (Fully Parameter Sharing), and FuPS+ID (FuPS with one-hot agent IDs). All policies use single-layer neural networks with controlled parameter counts (see Appendix I for details).
Figure 13: Mean evaluation reward (mean ± standard error) as a function of the number of agents/actions in the Specialisation (left) and Synchronisation (right) games. In the Specialisation game, vanilla policy gradients (PG, i.e. REINFORCE) with FuPS collapse as the team grows, whereas our identity-aware variant (PG-FuPS+ID) retains near-optimal performance. In the Synchronisation game, PG-NoPS performs well at small scales but degrades with more agents, while both PG-FuPS and PG-FuPS+ID remain at optimal reward across all scales.
Appendix F HyperMARL Details
F.1 HyperMARL Pseudocode
@@
4:   Agent embeddings $\{e^{i}\}_{i=1}^{n}$One-hot or orthogonally initialised learnable parameters}
@@
F.2 Variance of the HyperMARL Gradient Estimator
Unbiased estimator.
Following from Equation 4, we can write the unbiased estimator for HyperMARLs gradient as follows:
$\widehat{g}_{\mathrm{HM}}=\sum_{i=1}^{I}\underbrace{\nabla_{\!\psi}h_{\psi}^{\pi}(e^{i})}_{\mathbf{J}_{i}}\;\cdot\;\underbrace{\Bigl{(}\frac{1}{B}\sum_{b=1}^{B}\sum_{t=0}^{T-1}A(\mathbf{h}_{t}^{(b)},\mathbf{a}_{t}^{(b)})\,\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i,(b)}\!\mid h_{t}^{i,(b)})\Bigr{)}}_{\widehat{Z}_{i}}=\sum_{i=1}^{I}\mathbf{J}_{i}\,\widehat{Z}_{i},$
where $B$ trajectories $\{\tau^{(b)}\}_{b=1}^{B}$ are sampled i.i.d. and $\widehat{Z}_{i}$ is the empirical analogue of the observation-conditioned factor.
Assumptions.
(A1) trajectories are i.i.d.; (A2) all second moments are finite; (A3) $\psi,\theta,e^{i}$ are fixed during the backward pass.
Variance expansion.
Since each $\mathbf{J}_{i}$ is deterministic under (A3), we may factor them outside the covariance:
$\operatorname{Var}\bigl{(}\widehat{g}_{\text{\tiny HM}}\bigr{)}$ $=\operatorname{Cov}\!\Bigl{(}\sum_{i}\mathbf{J}_{i}\widehat{Z}_{i},\;\sum_{j}\mathbf{J}_{j}\widehat{Z}_{j}\Bigr{)}$ (by def. $\operatorname{Var}(X)=\operatorname{Cov}(X,X)$)
$=\sum_{i,j}\operatorname{Cov}\bigl{(}\mathbf{J}_{i}\widehat{Z}_{i},\;\mathbf{J}_{j}\widehat{Z}_{j}\bigr{)}$ (bilinearity of $\operatorname{Cov}$)
$=\sum_{i,j}\mathbf{J}_{i}\;\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})\;\mathbf{J}_{j}^{\!\top}$ (pull deterministic matrices out of $\operatorname{Cov}$)
Equation (LABEL:eq:variance_split_detail) makes explicit that all trajectory-induced covariance is captured $\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})$, while the agent-conditioned Jacobians $\mathbf{J}_{i}$ remain trajectory noise-free.
Mini-batch update and covariance.  Let $\widehat{Z}_{i}$ be the unbiased mini-batch estimate of $Z_{i}$ and $\widehat{g}_{\text{\tiny HM}}=\sum_{i}\mathbf{J}_{i}\widehat{Z}_{i}$ the stochastic update. Because every $\mathbf{J}_{i}$ is deterministic (wrt. to mini-batch),
$\operatorname{Var}\!\bigl{(}\widehat{g}_{\text{\tiny HM}}\bigr{)}=\sum_{i,j}\mathbf{J}_{i}\;\operatorname{Cov}(\widehat{Z}_{i},\widehat{Z}_{j})\;\mathbf{J}_{j}^{\!\top},$ (12)
(derivation in Appendix F.2). Equation (12) shows that HyperMARL first averages noise within each agent ($\widehat{Z}_{i}$) and only then applies $\mathbf{J}_{i}$. FuPS+ID, by contrast, updates the shared weights $\theta$ with every raw sample $A\,\nabla_{\theta}\log\pi_{\theta}[h,\mathrm{id}]$, leaving observation noise and agent ID entangled and making it susceptible to gradient interference [10, 49].
F.3 Scalability and Parameter Efficiency
@@
To reduce parameter count, techniques like shared hypernetworks, chunked hypernetworks [42, 9], or producing low-rank weight approximations, can be used. While naive implementations are parameter-intensive, this might be less critical in RL and MARL which commonly have smaller actor-critic networks. Moreover, HyperMARL’s near-constant scaling with agents suggests strong potential for large-scale MARL applications.
@@
F.4 Speed and Memory Usage
We examine the computational efficiency of HyperMARL compared to NoPS and FuPS by measuring inference speed (Figure 16(a)) and GPU memory usage (Figure 16(b)) as we scale the number of agents. The benchmarks were conducted using JAX on a single NVIDIA GPU (T4) with a recurrent (GRU-based) policy architecture. All experiments used fixed network sizes (64-dimensional embeddings and hidden layers) with a batch size of 128 and 64 parallel environments, allowing us to isolate the effects of varying agent count. Each measurement represents the average of 100 forward passes per configuration, with operations repeated across 10 independent trials.
(a) Forward pass inference time
(b) GPU memory usage
Figure 16: Computational efficiency scaling with number of agents. HyperMARL offers a balance between NoPS and FuPS. Notably, in real-world deployments, NoPS incurs additional data transfer and synchronisation costs not reflected here, further widening the efficiency gap.
The results demonstrate that HyperMARL offers a balance between the extremes represented by NoPS and FuPS. In practice, NoPS incurs additional data transfer and update costs, widening the efficiency gap.
Appendix G Experiment Details
G.1 Environments
@@
Dispersion (VMAS) [5]: A 2D environment where four agents collect unique food particles. This task requires specialised heterogeneous behaviours and resembles the Specialisation Game from Section E.1.
@@
Navigation (VMAS) [5]: Agents navigate in a 2D space to assigned goals, receiving dense rewards based on proximity. Configurations include shared goals (homogeneous), unique goals (heterogeneous), and mixed goals, where some agents share goals while others have unique ones.
@@
Multi-Agent MuJoCo (MAMuJoCo) [34]: A multi-agent extension of MuJoCo, where robot body parts (e.g., a cheetah’s legs) are modelled as different agents. Agents coordinate to perform efficient motion, requiring heterogeneous policies [49].
@@
SMAX (JaxMARL) [36]: Discrete action tasks with 2 to 20 agents on SMACv1- and SMACv2-style maps. FuPS baselines have been shown optimal for these settings [48, 16] indicating homogeneous behaviour is preferred here.
Table 4: Baseline Methods Selection and Justification. Selected methods (✓) were chosen based on their relevance to parameter sharing and specialisation/generalisation in MARL, while excluded methods (✗) did not align with our research objectives or had implementation constraints. Our experimental design systematically addresses key questions on agent specialisation and homogeneity, therefore we selected baselines with demonstrated strong performance in their respective settings, ensuring fair and rigorous comparison.
Method Category Selected Justification & Experimental Settings
IPPO [12]
(NoPS, FuPS+ID) NoPS/FuPS Established MARL baseline implementing both independent (NoPS) and fully shared (FuPS+ID) policy configurations. Tested in: Dispersion, Navigation, SMAX (two SMACv1 maps and two SMACv2 maps, with 10 and 20 agents).
MAPPO [48]
(NoPS, FuPS+ID) NoPS/FuPS Strong baseline with centralized critics for both NoPS and FuPS+ID architectures. Tested in: Dispersion, MAMuJoCo, SMAX (two SMACv1 maps and two SMACv2 maps, with 10 and 20 agents).
DiCo [6] Architectural Diversity Provides comparison with a method employing preset diversity levels that balances shared and non-shared parameters. Tested in: Dispersion and Navigation (as per original paper). Original hyperparameters used for $n=2$ agents; parameter sweeps conducted for $n>2$ to identify optimal diversity levels.
HAPPO [49] Sequential Updates Enables comparison with a method designed for heterogeneous behaviours using sequential policy updates with agent-specific parameters. Tested in: MAMuJoCo, selecting 4/6 scenarios from the original paper, including the challenging 17-agent humanoid task. Walker and Hopper variants were excluded as MAPPO and HAPPO performed similarly in these environments.
Kaleidoscope [26] Architectural Pruning Implemented for off-policy evaluation using its MATD3 implementation with tuned MaMuJoCo hyperparameters. Tested in: MAMuJoCo environments Ant-v2, HalfCheetah-v2, Walker2d-v2 (overlapping with our IPPO experiments), and Swimmer-v2-10x2 (highest agent count variant). Included to evaluate HyperMARL’s competitiveness against a method with ensemble critics and diversity loss, in an off-policy setting.
@@
SEAC [11] Shared Experience Focuses primarily on experience sharing rather than parameter sharing architecture, falling outside our research scope.
@@
SePS [10] Selective Parameter Sharing Requires pretraining phase, which extends beyond the scope of our current study focused on end-to-end learning approaches.
@@
CDAS [25] Intrinsic Reward Only implemented for off-policy methods and has been demonstrated to underperform FuPS/NoPS architectures [16], making it less suitable for our primary on-policy comparisons.
ROMA/RODE [43, 44] Role-based Shows limited practical performance advantages in comparative studies [11], suggesting other baselines provide more rigorous comparison points.
SNP-PS [23] Architectural Pruning No publicly available implementation, preventing direct, reproducible comparison.
G.2 HyperMARL Architecture Details
For the Dispersion and Navigations results (Sec. 5.2) we use feedforward architectures, where we use HyperMARL to generate both the actor and critic networks. For the MAPPO experiments in Section 5.2, for fairness in comparisons with HAPPO and MAPPO, we maintain the centralised critic conditioned on the global state and only use HyperMARL to generate the weights of the actors. For the recurrent IPPO experiments in Section 5.3, HyperMARL only generates the actor and critic feedforward weights, not the GRU weights.
G.2.1 Training and Evaluation